# SkyDiffusion: Street-to-Satellite Image Synthesis with Diffusion Models and BEV Paradigm

### v1

@@ 4.3 Comparison with State-of-the-Art Methods
On the suburban CVUSA and CVACT datasets, our SkyDiffusion method achieves the best results, with a 28.4% reduction in FID and a 17.4% improvement in SSIM compared to the state-of-the-art, demonstrating its superiority in synthesizing realistic and consistent and satellite images. The visual results in Fig. 4 show that GAN-based cross-view methods generate excessive artifacts and blurriness. While diffusion-based methods such as Aerial Diff [14] and Instr-p2p [1] produce highly realistic aerial views, they lack content correlation with the street-view images. In contrast, our method benefits from BEV perspective transformation, enabling effective capture of associative information. SkyDiffusion generates realistic images with consistent content layout, including road direction, tree placement, and building distribution.
Table 1: Quantitative comparison of different methods on CVUSA [42] and CVACT [18].
Method CVUSA CVACT FID ($\downarrow$) SSIM ($\uparrow$) PSNR ($\uparrow$) LPIPS ($\downarrow$) FID ($\downarrow$) SSIM ($\uparrow$) PSNR ($\uparrow$) LPIPS ($\downarrow$) X-Seq [24] 161.16 0.084 11.97 0.706 190.12 0.042 12.41 0.661 CDTE [38] 122.84 0.143 10.01 0.694 160.81 0.091 12.59 0.663 CUT [21] 72.52 0.122 12.12 0.695 62.34 0.101 13.02 0.662 Aerial Diff [14] 136.18 0.103 10.06 0.855 127.29 0.108 10.24 0.878 DreamBooth [27] 102.99 0.103 11.31 0.812 88.12 0.107 12.81 0.791 HawkI [15] 90.98 0.114 11.09 0.783 135.75 0.112 10.97 0.764 img2img-turbo [22] 77.95 0.103 12.64 0.685 73.24 0.103 12.19 0.679 Instr-p2p [1] 38.01 0.138 12.29 0.697 49.62 0.116 12.51 0.682 Ours 29.01 0.167 13.07 0.653 35.51 0.123 13.58 0.657
Figure 4: Qualitative comparison of synthesis results for suburban areas. Rows 1-3 show the results on the CVUSA dataset, and rows 4-6 show the results on the CVACT dataset.
Figure 5: Qualitative comparison of synthesis results for urban areas (VIGOR-Chicago).
Table 2: Quantitative comparison of different methods on VIGOR-Chicago.
Method Chicago FID ($\downarrow$) SSIM ($\uparrow$) PSNR ($\uparrow$) LPIPS ($\downarrow$) CUT [21] 69.63 0.171 13.82 0.666 Aerial Diff [14] 123.16 0.141 11.49 0.831 DreamBooth[27] 132.16 0.116 9.70 0.754 HawkI [15] 123.47 0.124 9.67 0.752 img2img-turbo [22] 80.10 0.135 11.28 0.690 Instr-p2p [1] 50.30 0.163 10.62 0.689 Ours(One-to-One) 46.35 0.185 10.48 0.665 Ours(Multi-to-One) 31.90 0.205 11.15 0.651
Table 3: Quantitative comparison of cross-dataset generalization test.
Method Cross-dataset FID ($\downarrow$) SSIM ($\uparrow$) PSNR ($\uparrow$) LPIPS ($\downarrow$) CVACT $\rightarrow$ VIGOR-Chicago CUT [21] 146.72 0.130 8.55 0.727 Ours (w/o light) 89.76 0.155 9.45 0.739 Ours 63.04 0.162 11.24 0.723 VIGOR-Chicago $\rightarrow$ CVACT CUT [21] 132.87 0.107 8.22 0.737 Ours (w/o light) 107.58 0.116 8.76 0.751 Ours 100.58 0.121 11.36 0.714
@@
In the urban VIGOR-Chicago dataset, SkyDiffusion achieved excellent performance with a 36.6% reduction in FID and a 20.5% improvement in SSIM compared to the state-of-the-art method. The visual results in Fig. 5 show that our method maintains consistent associative information such as crosswalks and lanes. While the non-associative information on building rooftops cannot be consistently generated, it still produces reasonable and realistic results.
Figure 6: Qualitative comparison of synthesis results for cross-dataset generalization.
4.4 Cross-Dataset Generalization
We trained the model on VIGOR [46] and tested it on the CVACT [18], and vice versa, to evaluate cross-dataset generation capability. Figure 6 shows the cross-dataset generation results of SkyDiffusion. The aim of this task is not to generate identical satellite images but to create scenes with consistent content layouts while exhibiting distinctly different styles. The visualization of the experimental results demonstrates that our method effectively preserves scene content such as road directions and intersections. Additionally, the images processed with the light manipulation module maintain the original content while better matching the illumination conditions of the target domain. Notably, we achieve this by using only a single random image from the training set of the target dataset to obtain the lighting conditions. However, due to the completely different architectural styles of the cities in the CVACT and VIGOR datasets, and because the models are trained on a single dataset, the generalization ability requires further improvement.
@@ 4.4 Cross-Dataset Generalization
Table 4: Ablation study of different modules on CVACT. "Street" represents directly using street-view image inputs, "Curved-BEV" denotes using Curved-BEV transformation, and "Light" stands for the Light Manipulation module.
Method Options CVACT Street Curved-BEV Light FID ($\downarrow$) SSIM ($\uparrow$) PSNR ($\uparrow$) LPIPS ($\downarrow$) CUT [21] 62.34 0.101 13.02 0.662 CUT [21] 51.71 0.121 13.14 0.658 Aerial Diff [14] 127.29 0.108 10.24 0.878 Aerial Diff [14] 83.34 0.114 11.15 0.831 Ours 62.21 0.115 11.95 0.682 Ours 63.60 0.118 13.49 0.681 Ours 36.48 0.117 13.19 0.642 Ours 35.51 0.123 13.58 0.657
@@
Figure 8: Qualitative analysis of ablation results for light manipulation.
4.5 Ablation Study
Curved-BEV controlled synthesis. We conducted ablation experiments on the CVACT dataset, with results shown in Table 4. By comparing the performance of different methods using Curved-BEV and street view inputs, we demonstrate that Curved-BEV consistently enhances performance across various architectures. Additionally, the proposed diffusion method achieves the best results. Further comparisons between existing BEV and Curved-BEV can be found in the supplementary materials.
Furthermore, the Multi-to-one method significantly improves metrics compared to the one-to-one mapping, demonstrating its effectiveness in dense urban scenes (Table 3). As shown in the visual results in Figure 7, using multiple street-view images for multi-to-one BEV mapping effectively addresses occlusions from tall objects like trees and buildings. This method effectively enhances the BEV sensing range in urban environments, facilitating the generation of large-scale satellite scenes.
Light manipulation. The ablation experiments in Table 4 and the cross-dataset generation results in Table 3 indicate that the Light Manipulation module aligns the lighting conditions of the synthesized images with those of the target domain images, significantly improving SSIM and PSNR metrics. The visualization results in Figure 8 further support this. This improvement is due to the consistent lighting conditions in the satellite data from CVACT and VIGOR-Chicago, which were captured in the same city. Using a small satellite image slice as a reference effectively captures the lighting conditions at the time of capture. Details on this module can be found in the supplementary materials.
5 Conclusion
In this study, we introduce SkyDiffusion, a novel approach specifically designed for street images to satellite images cross-view synthesis. The network operates solely with street images as input, utilizing a curved BEV method and diffusion models to generate realistic and content-consistent satellite images, similar to looking down from the sky. SkyDiffusion achieves state-of-the-art performance in both content consistency and image realism on suburban (CVUSA & CVACT) and urban (VIGOR-Chicago) cross-view datasets, demonstrating significant potential in real-world image synthesis in complex application scenes.

### v2

@@ Methods
On the suburban CVUSA and CVACT datasets, our SkyDiffusion method achieved the outstanding results, as shown in Table 1. Compared to state-of-the-art methods, it reduced FID by 25.83% and increased SSIM by 13.89%, demonstrating its superiority in synthesizing realistic and consistent satellite images. The visual results in Fig. 4 show that GAN-based cross-view methods generate excessive artifacts and blurriness. While diffusion-based methods such as Aerial Diff and Instr-p2p produce highly realistic aerial views, they lack content correlation with the street-view images. In contrast, our method benefits from BEV perspective transformation, enabling effective capture of cross-view associative information. SkyDiffusion generates realistic images with consistent content layout, including road direction, tree placement, and building distribution. Additional visual results can be found in the supplementary materials.
@@
In the urban VIGOR-Chicago dataset, SkyDiffusion reduced FID by 9.96% and improved SSIM by 14.11% compared to the state-of-the-art method, as shown in Table 1. For fair comparison, the VIGOR results for SkyDiffusion use only a single street-view image. The results using Multi-to-One BEV generation with multiple street-view images can be found in the ablation study. The visual results in Fig. 4 show that our method maintains consistent associative information such as crosswalks and lanes. While the non-associative information on building rooftops cannot be consistently generated, it still produces reasonable and realistic results.
Ablation Study
Curved-BEV controlled synthesis. We conducted ablation experiments on the three datasets, with results shown in Table 2. Compared to directly using street-view images as input, the Curved-BEV method improves performance across multiple metrics by transforming street-view images into satellite views for domain alignment. This indicates that the Curved-BEV method aids in synthesizing more content-consistent satellite images. Furthermore, the Multi-to-one method further improves metrics compared to the one-to-one mapping, demonstrating its effectiveness in dense urban scenes. As shown in the visual results in Figure 5, using multiple street-view images for multi-to-one BEV mapping effectively addresses occlusions from tall objects like trees and buildings. This method effectively enhances the BEV sensing range in urban environments, facilitating the generation of large-scale satellite scenes. Additional ablation study results can be found in the supplementary materials.
Method FID ($\downarrow$) SSIM ($\uparrow$) PSNR ($\uparrow$) LPIPS ($\downarrow$)
CVUSA
Baseline 32.45 0.149 12.63 0.650
Ours (C-BEV) 29.18 0.168 14.58 0.635
CVACT
Baseline 62.21 0.115 11.95 0.682
Ours (C-BEV) 35.51 0.123 13.58 0.657
VIGOR-Chicago
Baseline 53.27 0.170 10.38 0.666
Ours (C-BEV) 45.29 0.186 11.69 0.661
Ours (C-BEV Multi) 31.90 0.205 11.15 0.651
@@
Table 2: Ablation study of the Curved-BEV module. "Baseline" represents directly using street-view image,"C-BEV" denotes using Curved-BEV transformation, and "Multi" stands for Multi-to-One strategy.
@@
Light manipulation. The ablation experiments in Table 3 indicate that the Light Manipulation module aligns the lighting conditions of the synthesized images with those of the target domain images, improving SSIM and PSNR metrics. This is because satellites typically capture large areas with relatively uniform overall brightness. Using a small satellite image slice as a reference effectively captures the lighting conditions at the time of capture. Due to the dispersed distribution of the CVUSA data and the more complex lighting conditions, the improvements are not as pronounced as on other datasets. The light processing module maintains the original image content while providing more flexible lighting conditions for the synthesized satellite images. More visualization results can be found in the supplementary material.
Method FID ($\downarrow$) SSIM ($\uparrow$) PSNR ($\uparrow$) LPIPS ($\downarrow$)
CVUSA
Ours (w/o light) 29.02 0.167 14.07 0.643
Ours 29.18 0.168 14.58 0.635
CVACT
Ours (w/o light) 36.48 0.117 12.85 0.665
Ours 35.51 0.123 13.58 0.657
VIGOR-Chicago
Ours (w/o light) 46.77 0.179 10.46 0.666
Ours 45.29 0.186 11.69 0.661
Table 3: Ablation study of the light manipulation module.
Figure 6: Synthesis results for cross-dataset generalization.
Method FID ($\downarrow$) SSIM ($\uparrow$) PSNR ($\uparrow$) LPIPS ($\downarrow$)
CVACT $\rightarrow$ VIGOR-Chicago
Instr-p2p 100.80 0.152 9.14 0.717
Ours (w/o light) 89.76 0.155 9.45 0.739
Ours 63.04 0.162 11.24 0.723
VIGOR-Chicago $\rightarrow$ CVACT
Instr-p2p 121.29 0.113 8.90 0.759
Ours (w/o light) 107.58 0.116 8.76 0.751
Ours 100.58 0.121 11.36 0.714
Table 4: Cross-dataset generalization assessment.
Cross-Dataset Generalization
We trained the model on CVACT and tested it on the VIGOR-Chicago, and vice versa, to evaluate cross-dataset generation capability, as shown in Table 4. Compared to Instruct pix2pix, our method (w/o light) demonstrates superior performance across metrics. Figure 6 shows the synthesis results for cross-dataset generalization. Our method effectively preserves scene content such as road directions and intersections. However, due to the distinct city styles of the CVACT and VIGOR datasets and the fact that the model was trained on a single dataset, there are noticeable differences in elements such as buildings and lighting conditions. The aim of this task is not to generate identical satellite images but to create scenes with consistent content layouts while exhibiting distinctly different styles. Additionally, we can flexibly use the light manipulation module to obtain light information from a random image in the target dataset’s training set, better matching the target domain’s lighting conditions.
Conclusion
In this study, we introduce SkyDiffusion, a novel approach specifically designed for street images to satellite images cross-view synthesis. SkyDiffusion operates solely with street images as input, utilizing a BEV Paradigm and diffusion models to generate satellite images. SkyDiffusion achieves state-of-the-art performance in both content consistency and image realism on suburban and urban cross-view datasets, demonstrating its superior capabilities in cross-view synthesis tasks. The satellite and street-view domains bridging method, Curved-BEV, within this approach also holds valuable insights for tasks such as cross-view retrieval and geolocation. In the future work, we plan to extend our method to more cities and further refine it for specific downstream applications, such as urban planning and virtual tourism.